# Continuous fields of interval algebras

### v1

@@ 1. Introduction
This work addresses the following question: Can we classify continuous fields of unital $\operatorname{AI}$-algebras? While we aim for broad generality, we acknowledge the presence of natural obstructions and consequent restrictions. E.g. on the dimension of the topological ‘base’ space or on the global behavior of the fibers will appear. We first thoroughly establish the main definitions and properties of continuous fields of $C^{*}$-algebras and their morphisms. We then employ a local-to-global strategy. First, we extend the existence part of Robert’s classification to morphisms of continuous fields using an elegant version of Michael’s selection theorem. Second, we extend the uniqueness part. using a technical gluing process. Finally, we introduce the concept of continuous fields of Cuntz semigroups, leading to a functorial classification of certain continuous fields of $\operatorname{AI}$-algebras. The main results of this manuscript are detailed below.
@@
Then there exists a continuous field morphism $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$, unique up to approximate unitary equivalence, such that $\operatorname{Cu}(\phi)=\alpha$.
@@
Acknowledgments. The author would like to thank S. Bhattacharjee for introducing him to the world of continuous fields of $C^{*}$-algebras and M. Dădărlat for fruitful discussions.
@@ 2. An intro on continuous fields of $C^{*}$-algebras
Recall that by a section, we mean a map $a\colon X\rightarrow\bigsqcup_{x}A_{x}$ such that $a(x)\in A_{x}$ for all $x\in X$. Equivalently, a section is an element of $\prod_{x}A_{x}$.
[…]
@@
Note that the above definition has a ‘$C^{*}$-algebraic flavor’ as opposed to a more traditional picture which has a more ‘topological flavor’. Namely, continuous fields are sometimes defined as a triple $(E,X,\pi\colon\mathcal{A}\rightarrow X)$ where $E$ and $X$ are topological spaces (respectively, the global and the base space) and $\pi$ is a continuous surjection satisfying specific properties. In this picture, the fibers would consist of the family $\{\pi^{-1}(\{x\})\}_{x}$. See [4] for more on this.
@@
$V_{\epsilon}^{c}:=\{a_{v}\in A_{v}\mid v\in V\text{ and }\|a_{v}-c(v)\|<\epsilon\}$
where $V\subseteq X$ is an open set, $c\in\mathcal{A}$ and $\epsilon>0$.
@@
[…]
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ be a continuous field of $C^{*}$-algebras and let $(a_{x})_{x}$ be a section. Then the following are equivalent.
(i) $(a_{x})_{x}\in\mathcal{A}$.
(ii) For any $x\in X$ and any $\epsilon>0$, there exist an open neighborhood $V$ of $x$ and some $c\in\mathcal{A}$ such that $a_{v}\in V_{\epsilon}^{c}$ for all $v\in V$.
(iii) The map $x\mapsto a_{x}$ is continuous with respect to $\tau_{\mathcal{A}}$.
@@
Let us write $f\colon x\mapsto a_{x}$ during the proof. The implication (i) $\Rightarrow$ (ii) is obvious.
Assume that $(a_{x})_{x}$ satisfies (ii). Consider a basic open set $V_{\epsilon}^{c}$. Let $v\in f^{-1}(V_{\epsilon}^{c})$ and write $C=\|a_{v}-c(v)\|<\epsilon$. Fix $\delta>0$ such that $3\delta+C<\epsilon$. Using the assumption, we know that there exists $d\in\mathcal{A}$ and an open neighborhood $W$ of $v$ such that $\|a_{w}-d(w)\|<\delta$ for any $w\in W$. We deduce that $\|(c-d)(v)\|\leq C+\delta$. By the continuity of $\nu_{(c-d)}$, we can find an open neighborhood $Z$ of $v$ such that $Z\subseteq W\cap f^{-1}(V_{\epsilon}^{c})$ and such that $\|(c-d)(z)\|\leq C+2\delta$ for all $z\in Z$. Finally, we compute that $\|a_{z}-c(z)\|\leq\delta+C+2\delta<\epsilon$. As a result, we see that $Z\subseteq f^{-1}(V_{\epsilon}^{c})$ and we conclude that $(a_{x})_{x}$ satisfies (iii).
Assume that $(a_{x})_{x}$ satisfies (iii). Let $x\in X$ and $\epsilon>0$. By the surjectivity of $\pi_{x}\colon\mathcal{A}\rightarrow A_{x}$, we can find some $c\in\mathcal{A}$ such that $c(x)=a_{x}$. Consider $V:=f^{-1}(X_{\epsilon}^{c})$. By continuity of $f$, $V$ is an open neighborhood of $x$ and hence, $a_{x}\in V_{\epsilon}^{c}$. This shows that $(a_{x})_{x}$ satisfies (ii).
Assume that $(a_{x})_{x}$ satisfies (ii), again. Let $\epsilon>0$. By compactness of $X$, there exist a finite open cover $\{V_{k}\}_{1}^{l}$ of $X$ together with finitely many elements $\{c_{k}\}_{1}^{l}$ of $\mathcal{A}$ such that for any $v\in X$, we have $a_{v}\in(V_{k})_{\epsilon}^{c_{k}}$ for some $k$. Consider a partition $\{\lambda_{k}\}_{1}^{l}$ subordinated to the open cover $\{V_{k}\}_{1}^{l}$ and construct $c_{\epsilon}:=\sum_{1}^{l}\lambda_{k}c_{k}$. We have that $c_{\epsilon}\in\mathcal{A}$ and that $\|c_{\epsilon}(x)-a_{x}\|<\epsilon$ for all $x\in X$. A standard completeness argument yields (i). ∎
@@
We now define morphisms between continuous fields of $C^{*}$-algebras, again following the ‘$C^{*}$-algebraic flavor’.
[…]
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be a continuous field of $C^{*}$-algebras over the same space $X$. A continuous field morphism is a *-homomorphism $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$ which induces a family of *-homomorphisms $\{\phi_{x}\colon A_{x}\rightarrow B_{x}\}_{x}$ such that the following diagram commutes for all $x\in X$
@@
The following proposition shows a couple statements that are respectively sufficient and necessary for a field of *-homomorphisms to be continuous field morphism, which both turn out to be equivalent. Finally, we mention that these morphisms induce continuous maps between the global spaces of the continuous fields at hand.
[…]
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be a continuous field of $C^{*}$-algebras. Let $(\phi_{x}\colon A_{x}\rightarrow B_{x})_{x}$ be a family of *-homomorphisms. The following are equivalent.
@@
(iv) The map $\tilde{\phi}\colon\bigsqcup_{x}A_{x}\rightarrow\bigsqcup_{x}B_{x}$ sending $a_{x}\mapsto\phi_{x}(a(x))$ is $(\tau_{\mathcal{A}},\tau_{\mathcal{B}})$-continuous.
@@
Let us check that (ii) implies (iv). Let $V_{\epsilon}^{b}$ be a basic open set of $\tau_{\mathcal{B}}$ and let $a_{v}\in\tilde{\phi}^{-1}(V_{\epsilon}^{b})$. We know that $\|\phi_{v}(a_{v})-b(v)\|<\epsilon$. By the surjectivity of $\pi_{v}\colon\mathcal{A}\rightarrow A_{v}$, we can find some $a\in\mathcal{A}$ such that $a(v)=a_{v}$. By assumption $c:=(\phi_{x}(a(x)))_{x}$ is an element of $\mathcal{B}$. By the continuity of $\nu_{c-b}$, we conclude that there exists an open neighborhood $W$ of $v$ such that $W_{\delta}^{a}\subseteq\tilde{\phi}^{-1}(V_{\epsilon}^{b})$ for some $\delta>0$ small enough, from which the result follows. ∎
@@ 3. From local to global
Based on the work of several, it is known that the Cuntz semigroup classifies *-homomorphisms from a class of unital $C^{*}$-algebras with trivial $\mathrm{K}_{1}$-groups to stable rank one $C^{*}$-algebras. Let us recall a weaker result that will suffice our needs.
[…]
Let $A,B$ be unital separable $\operatorname{AI}$-algebras. Let $\alpha\colon\operatorname{Cu}(A)\rightarrow\operatorname{Cu}(B)$ be a $\operatorname{Cu}$-morphism satisfying $\alpha([1_{A}])\leq[1_{B}]$.
@@ A. The existence
A multi-valued function is an assignment $\Theta\colon X\rightarrow 2^{Y}$ where $X,Y$ are topological spaces and $2^{Y}$ denotes the set of all (non-empty) subsets of $Y$. The function $\Theta$ is said to be lower-semicontinuous if for any open set $V\subseteq Y$, the set $\{x\in X\mid\Theta(x)\cap V\}$ is open in $X$. A continuous selection for $\Theta$ is a continuous map $\phi\colon X\rightarrow Y$ such that $f(x)\in\Theta(x)$ for any $x\in X$.
Below, we state a weaker version of the selection theorem that will suffice our needs. Let us mention that a similar version of the theorem has been used in the work of [11].
[…]
[13, Theorem 1.2] Let $X$ be a compact Hausdorff space of dimension (at most) one. Let $(Y,d)$ be a metric space and let $\mathcal{S}$ be a family of closed non-empty subsets of $Y$ satisfying the following.
@@
Let us outline how we will invoke Michael’s selection theorem in our context, yielding to the existence of a continuous field morphism $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$ lifting a given $\operatorname{Cu}$-morphism $\alpha\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(\mathcal{B})$ descending to the fibers.
Firstly, we define a topology $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$ on the global space $\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$, which induces the point-norm topology on each $\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$. We then show that this topology is metrizable under suitable separability conditions of the fibers. (The metric space obtained will play the role of $Y$ in the selection theorem.) Subsequently, under a suitable projective property, we exhibit a continuity criterion for elements in $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$, based on the topology $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$, similarly as done in Section 2 for elements in $\mathcal{A}$.
Secondly, we consider $\mathcal{S}(x):=\{\phi_{x}\in\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\mid\operatorname{Cu}(\phi_{x})=\alpha_{x}\}$ for any $x\in X$. By Robert’s result, we deduce that these sets are non-empty and complete. Path-connectedness will follow from a result of [5].
Lastly, under the uniform path property, we apply the selection theorem to conclude via the continuity criterion. We end the section with a discussion on specific cases where the uniform path property and the projective property are automatically satisfied.
@@
If moreover, we assume that $A$ is a separable $C^{*}$-algebra, the space $(\operatorname{Hom}_{C^{*}}(A,B),\tau_{pn})$ becomes metrizable as follows. Let $(a_{k})_{k}$ be a dense sequence of self-adjoint elements of $A$. For any $\phi,\psi\in\operatorname{Hom}_{C^{*}}(A,B)$, we set
$d(\phi,\psi)=\sum_{0}^{\infty}\frac{\|\phi(a_{k})-\psi(a_{k})\|}{2^{k}\|a_{k}\|}$
It can be checked that $d$ is a well-defined metric on $\operatorname{Hom}_{C^{*}}(A,B)$.
[…]
The space $(\operatorname{Hom}_{C^{*}}(A,B),\tau_{pn})$ is completely metrizable whenever $A$ is a separable $C^{*}$-algebra.
More precisely, any of the above metrics $d$ induces $\tau_{pn}$ and turns $(\operatorname{Hom}_{C^{*}}(A,B),d)$ into a complete metric space.
@@
The fact that the metric $d$ is well-defined and induces $\tau_{pn}$ can be checked via standard arguments and is left to the reader. Regarding completeness, we observe that any $d$-Cauchy sequence $(\phi_{n})_{n}$ induces a .$\|.\|_{B}$-Cauchy sequence $(\phi_{n}(a_{k}))_{n}$ for any self-adjoint $a_{k}\in A$ involved in the construction of $d$. As a consequence, we define $\phi\colon A\rightarrow B$ via the assignment $a_{k}\mapsto\lim_{n}(\phi_{n}(a_{k}))$ that we have extended continuously from the dense sequence $(a_{k})_{k}$ to $A$. It is readily seen that $\phi$ is a linear map preserving the *-operation. Finally, for any $\epsilon>0$, any $k,l\in\mathbb{N}$, we can find $m_{\epsilon}$ such $\|\phi(a_{k})\phi(a_{l})-\phi_{m_{\epsilon}}(a_{k})\phi_{m_{\epsilon}}(a_{l})\|<\epsilon/2$ and $\|\phi_{m_{\epsilon}}(a_{k}a_{l})-\phi(a_{k}a_{l})\|<\epsilon/2$ which yields that $\phi$ is multiplicative on the dense subset $\{a_{k}\}_{k}$. We conclude that $\phi$ is a *-homomorphism by a standard continuity argument. ∎
@@
$V_{(F,\epsilon)}^{\chi}:=\{\phi_{v}\in\operatorname{Hom}_{C^{*}}(A_{v},B_{v})\mid v\in V\text{ and }\phi_{v}\simeq_{(F,\epsilon)}\chi_{v}\}$
@@
This topology will allow us to link elements of $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ with $(\tau_{(\mathcal{A}\rightarrow\mathcal{B})})$-continuous sections of the canonical surjection $\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\relbar\joinrel\twoheadrightarrow X$. In the process, we are obtaining a continuity criterion alike to the one of Section 2. First, we need the followings.
[…]
@@
Assume that $\phi_{x}\simeq_{(F_{x},\epsilon)}\psi_{x}$ for some $x\in X$. Then there exists an open neighborhood $V$ of $x$ such that $\phi_{v}\simeq_{(F_{v},\epsilon)}\psi_{v}$ for all $v\in V$.
@@
We are now ready for the continuity criterion for continuous field morphisms.
[…]
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be continuous fields of $C^{*}$-algebras. Let $(\phi_{x})_{x}$ be a section of *-homomorphisms. Consider the following statements.
(i) $(\phi_{x})_{x}\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$.
(ii) For any $x\in X$, any $F\subseteq\mathcal{A}$ finite and any $\epsilon>0$, there exist an open neighborhood $V$ of $x$ and some $\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ such that $\phi_{v}\in V_{(F,\epsilon)}^{\chi}$ for all $v\in V$.
(iii) $x\mapsto\phi_{x}$ is continuous with respect to $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$.
Then (i) and (ii) are equivalent and imply (iii). The converse holds whenever the set $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ satisfies (PP).

### v2

@@ 1. Introduction
This work addresses the following question: Can we classify continuous fields of unital $\operatorname{AI}$-algebras? While we aim for broad generality, we acknowledge the presence of natural obstructions and consequent restrictions. E.g. on the dimension of the topological ‘base’ space or on the global behavior of the fibers. After thoroughly establish the main definitions and properties of continuous fields of $C^{*}$-algebras and their morphisms, we employ a local-to-global strategy. Firstly, we extend the existence part of Robert’s classification to morphisms of continuous fields using an elegant version of Michael’s selection theorem. Secondly, we extend the uniqueness part using a technical gluing process. Finally, we introduce the concept of continuous fields of Cuntz semigroups, leading to a functorial classification of certain continuous fields of $\operatorname{AI}$-algebras. We expose our main result below.
@@
Let $\mathcal{A},\mathcal{B}\in\mathcal{C}$. Let $\alpha\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(\mathcal{B})$ be any scaled $\operatorname{Cu}$-morphism descending to the fibers. Then there exists a continuous field morphism $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$, unique up to approximate unitary equivalence, such that $\operatorname{Cu}(\phi)=\alpha$.
@@
Acknowledgments. The author would like to thank S. Bhattacharjee for introducing him to the world of continuous fields of $C^{*}$-algebras and M. Dădărlat for sharing an enlightening example. Also, the referee for making pertinent comments that have improved the exposition of the manuscript.
@@ 2. An intro on continuous fields of $C^{*}$-algebras
Recall that by a section, we mean a map $a\colon X\rightarrow\bigsqcup_{x}A_{x}$ such that $a(x)\in A_{x}$, for all $x\in X$. (Here and in the rest of this work, we freely identify $A_{x}$ with its image in $\bigsqcup_{x}A_{x}$.) Equivalently, a section is an element of $\prod_{x}A_{x}$.
Definition 2.1.
@@
Note that the above definition has a $C^{*}$-algebraic flavor’ as opposed to a more traditional picture which has a more ‘topological flavor’. Namely, continuous fields are sometimes defined as a triple $(E,X,\pi\colon\mathcal{A}\rightarrow X)$ where $E$ and $X$ are topological spaces (respectively, the global and the base space) and $\pi$ is a continuous surjection satisfying specific properties. In this picture, the fibers would consist of the family $\{\pi^{-1}(\{x\})\}_{x}$. See [5] for more on this.
@@
$V_{\epsilon}^{c}:=\{(\xi,x)\in\bigsqcup_{x\in X}A_{x}\mid x\in V\text{ and }\|\xi-c(x)\|<\epsilon\}$
where $V\subseteq X$ is an open set, $c\in\mathcal{A}$ and $\epsilon>0$. (Again, we have freely identified $A_{x}$ with its image in $\bigsqcup_{x}A_{x}$, and the above norm is computed in $A_{x}$.)
@@
Proposition 2.2 (Continuity Criterion).
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ be a continuous field of $C^{*}$-algebras and let $a\colon X\rightarrow\bigsqcup_{x}A_{x}$ be a section. Then the following are equivalent.
(i) $a\in\mathcal{A}$.
(ii) For any $x\in X$ and any $\epsilon>0$, there exist an open neighborhood $V$ of $x$ and some $c\in\mathcal{A}$ such that $a(v)\in V_{\epsilon}^{c}$, for all $v\in V$.
(iii) $a$ is continuous with respect to $\tau_{\mathcal{A}}$.
@@
The implication (i) $\Rightarrow$ (ii) is obvious.
Assume that $a$ satisfies (ii). Consider a basic open set $V_{\epsilon}^{c}$ such that $a^{-1}(V_{\epsilon}^{c})\neq\emptyset$. Let $v\in a^{-1}(V_{\epsilon}^{c})$ and write $C=\|a(v)-c(v)\|<\epsilon$. Fix $\delta>0$ such that $3\delta+C<\epsilon$. Using the assumption, we know that there exists $d\in\mathcal{A}$ and an open neighborhood $W$ of $v$ such that $\|a(w)-d(w)\|<\delta$, for any $w\in W$. We deduce that $\|(c-d)(v)\|\leq C+\delta$. By the continuity of $\nu_{(c-d)}$, we can find an open neighborhood $Z$ of $v$ such that $\|(c-d)(z)\|\leq C+2\delta$ for all $z\in Z$. Finally, we compute that $\|a(y)-c(y)\|\leq\delta+C+2\delta<\epsilon$, for all $y\in Z\cap W$. As a result, we see that $Z\cap W\subseteq a^{-1}(V_{\epsilon}^{c})$ and we conclude that $a$ satisfies (iii).
Assume that $a$ satisfies (iii). Let $x\in X$ and $\epsilon>0$. By the surjectivity of $\pi_{x}\colon\mathcal{A}\rightarrow A_{x}$, we can find some $c\in\mathcal{A}$ such that $c(x)=a(x)$. By the continuity of $a$, we deduce that $V:=a^{-1}(X_{\epsilon}^{c})$ is an open neighborhood of $x$ such that $a(v)\in V^{c}_{\epsilon}$ for all $v\in V$, yielding (ii).
Assume that $a$ satisfies (ii), again. Let $\epsilon>0$. By compactness of $X$, there exist a finite open cover $\{V_{k}\}_{1}^{l}$ of $X$ together with finitely many elements $\{c_{k}\}_{1}^{l}$ of $\mathcal{A}$ such that for any $v\in X$, we have $a(v)\in(V_{k})_{\epsilon}^{c_{k}}$ for some $k$. Consider a partition $\{\lambda_{k}\}_{1}^{l}$ subordinated to the open cover $\{V_{k}\}_{1}^{l}$ and construct $c_{\epsilon}:=\sum_{1}^{l}\lambda_{k}c_{k}$. We have that $c_{\epsilon}\in\mathcal{A}$ and that $\|c_{\epsilon}(x)-a(x)\|<\epsilon$ for all $x\in X$. A standard completeness argument yields (i). ∎
@@
We now define morphisms between continuous fields of $C^{*}$-algebras, again following the $C^{*}$-algebraic flavor’.
Definition 2.4.
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be continuous fields of $C^{*}$-algebras over the same space $X$. A continuous field morphism is a *-homomorphism $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$ which induces a family of *-homomorphisms $\{\phi_{x}\colon A_{x}\rightarrow B_{x}\}_{x}$ such that the following diagram commutes for all $x\in X$
@@
The following proposition shows a couple of statements that are respectively sufficient and necessary for a field of *-homomorphisms to be a continuous field morphism, which both turn out to be equivalent. Finally, we mention that these morphisms induce continuous maps between the global spaces of the continuous fields at hand.
Proposition 2.5.
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be continuous fields of $C^{*}$-algebras. Let $(\phi_{x}\colon A_{x}\rightarrow B_{x})_{x}$ be a family of *-homomorphisms. The following are equivalent.
@@
(iv) The map $\tilde{\phi}\colon\bigsqcup_{x}A_{x}\rightarrow\bigsqcup_{x}B_{x}$ sending $(\xi,x)\mapsto(\phi_{x}(\xi),x)$ is $(\tau_{\mathcal{A}},\tau_{\mathcal{B}})$-continuous.
@@
Let us check that (ii) implies (iv). Let $V_{\epsilon}^{b}$ be a basic open set of $\tau_{\mathcal{B}}$ and let $(\xi,v)\in\tilde{\phi}^{-1}(V_{\epsilon}^{b})$. We know that $\|\phi_{v}(\xi)-b(v)\|<\epsilon$. By the surjectivity of $\pi_{v}\colon\mathcal{A}\rightarrow A_{v}$, we can find some $a\in\mathcal{A}$ such that $a(v)=\xi$. By assumption $c:=(\phi_{x}(a(x)))_{x}$ is an element of $\mathcal{B}$. By the continuity of $\nu_{c-b}$, we conclude that there exists an open neighborhood $W$ of $v$ such that $W_{\delta}^{a}\subseteq\tilde{\phi}^{-1}(V_{\epsilon}^{b})$ for some $\delta>0$ small enough, from which the result follows. ∎
@@ 3. From local to global
Based on the works of Ciuperca, Elliott, Robert and Santiago, it is known the Cuntz semigroup classifies *-homomorphisms from a class of unital $C^{*}$-algebras with trivial $\mathrm{K}_{1}$-groups to stable rank one $C^{*}$-algebras. See [12, 21, 22]. Let us recall a weaker version that will suffice our needs.
Theorem 3.1.
([22, Theorem 1.0.1]) Let $A,B$ be unital separable $\operatorname{AI}$-algebras and let $\alpha\colon\operatorname{Cu}(A)\rightarrow\operatorname{Cu}(B)$ be a $\operatorname{Cu}$-morphism satisfying $\alpha([1_{A}])\leq[1_{B}]$.
@@ A. The existence
Recall that a multi-valued function is an assignment $\Theta\colon X\rightarrow 2^{Y}$ where $X,Y$ are topological spaces and $2^{Y}$ denotes the set of all (non-empty) subsets of $Y$. A continuous selection for $\Theta$ is a continuous map $\phi\colon X\rightarrow Y$ such that $f(x)\in\Theta(x)$ for any $x\in X$. Finally, the multi-valued function $\Theta$ is said to be lower-semicontinuous, if for any open set $V\subseteq Y$, the set $\{x\in X\mid\Theta(x)\cap V\neq\emptyset\}$ is open in $X$.
Below is a weaker version of the selection theorem that will suffice our needs. Let us mention that a similar version of the theorem has been used in [16].
Theorem 3.3.
[18, Theorem 1.2] Let $X$ be a compact Hausdorff space of dimension (at most) one. Let $(Y,d)$ be a metric space and let $\mathcal{S}$ be a family of closed non-empty subsets of $Y$ satisfying the following.
@@
Let us outline how we will invoke Michael’s selection theorem in our context, yielding the existence of a continuous field morphism $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$ lifting a given $\operatorname{Cu}$-morphism $\alpha\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(\mathcal{B})$ descending to the fibers.
Firstly, we define a topology $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$ on the global space $\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$, which induces the point-norm topology on each ‘fiber’ $\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$. We then show that this topology is metrizable under suitable separability conditions of the ‘fibers’ $A_{x}$. (The metric space obtained will play the role of $Y$ in the selection theorem.) Subsequently, under a suitable projective property, we exhibit a continuity criterion for elements in $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$, based on the topology $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$, similarly as done in Section 2 for elements of $\mathcal{A}$.
Secondly, we consider $\mathcal{S}_{\alpha}(x):=\{\phi_{x}\in\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\mid\operatorname{Cu}(\phi_{x})=\alpha_{x}\}$ for any $x\in X$. By Robert’s result, we deduce that these sets are non-empty and complete. Path-connectedness will follow from a result of [6].
Lastly, under assumptions on the continuous fields at hand, we show the uniform path property of $\mathcal{S}_{\alpha}$ and the lower-semicontinuity of the map sending $x\mapsto\mathcal{S}_{\alpha}(x)$. We finally apply the selection theorem to conclude our existence result via the continuity criterion.
We end the section with a discussion on specific cases where the uniform path property and the projective property are automatically satisfied.
@@
If moreover $A$ is a separable $C^{*}$-algebra, the space $(\operatorname{Hom}_{C^{*}}(A,B),\tau_{pn})$ (where $\tau_{pn}$ denotes the point-norm topology as defined above) becomes completely metrizable. We give below an explicit construction of such a complete metric, for the reader’s convenience.
Let $(a_{k})_{k\in\mathbb{N}}$ be a dense subset of self-adjoint elements of $A$. For any $\phi,\psi\colon A\rightarrow B$, we set
$d(\phi,\psi)=\sum_{k=0}^{\infty}\frac{\|\phi(a_{k})-\psi(a_{k})\|}{2^{k}\|a_{k}\|}$
Proposition 3.4.
The above metric $d$ induces the point-norm topology $\tau_{pn}$ and turns $(\operatorname{Hom}_{C^{*}}(A,B),d)$ into a complete metric space.
@@
The fact that the metric $d$ is well-defined and induces $\tau_{pn}$ can be checked via standard arguments and is left to the reader. Regarding completeness, we observe that any $d$-Cauchy sequence $(\phi_{n})_{n}$ induces a $\|.\|_{B}$-Cauchy sequence $(\phi_{n}(a_{k}))_{n}$ for any self-adjoint $a_{k}\in A$ involved in the construction of $d$. As a consequence, we define $\phi\colon A\rightarrow B$ via the assignment $a_{k}\mapsto\lim_{n}(\phi_{n}(a_{k}))$ that we have extended continuously from the dense sequence $(a_{k})_{k}$ to $A$. It is readily seen that $\phi$ is a linear map preserving the *-operation. Finally, for any $\epsilon>0$, any $k,l\in\mathbb{N}$, we can find $m_{\epsilon}$ such $\|\phi(a_{k})\phi(a_{l})-\phi_{m_{\epsilon}}(a_{k})\phi_{m_{\epsilon}}(a_{l})\|<\epsilon/2$ and $\|\phi_{m_{\epsilon}}(a_{k}a_{l})-\phi(a_{k}a_{l})\|<\epsilon/2$ which yields that $\phi$ is multiplicative on the dense subset $\{a_{k}\}_{k}$. We conclude that $\phi$ is a *-homomorphism by a standard continuity argument. ∎
@@
$V_{(F,\epsilon)}^{\chi}:=\{(\xi,x)\in\bigsqcup_{x\in X}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\mid x\in V\text{ and }\xi\simeq_{(F,\epsilon)}\chi_{x}\}$
@@
This topology will allow us to link elements of $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ with $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$-continuous sections of the canonical surjection $\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\relbar\joinrel\twoheadrightarrow X$. In the process, we are obtaining a continuity criterion alike to the one of Section 2. First, we need the following.
Definition 3.5.
@@
Assume that $\phi_{x}\simeq_{(F_{x},\epsilon)}\psi_{x}$, for some $x\in X$. Then there exists an open neighborhood $V$ of $x$ such that $\phi_{v}\simeq_{(F_{v},\epsilon)}\psi_{v}$, for all $v\in V$.
@@
We are now ready for the continuity criterion for continuous field morphisms. (Again, we freely identify $\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$ with its image in $\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$.)
Proposition 3.7 (Continuity Criterion).
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be continuous fields of $C^{*}$-algebras. Let $(\phi_{x})_{x}$ be a family of *-homomorphisms. Consider the following statements.
(i) The map $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$ sending $a\rightarrow(x\mapsto\phi_{x}(a(x)))$ is a well-defined continuous field morphism.
(ii) For any $x\in X$, any $F\subseteq\mathcal{A}$ finite and any $\epsilon>0$, there exist an open neighborhood $V$ of $x$ and some $\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ such that $\phi_{v}\in V_{(F,\epsilon)}^{\chi}$, for all $v\in V$.
(iii) The map from $X\rightarrow\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$ sending $x\mapsto\phi_{x}$ is continuous with respect to $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$.
Then (i) and (ii) are equivalent and imply (iii). The converse holds whenever the set $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ satisfies $\rm(PP)$.